Circular No. 5817

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 KA_{2}

An asteroidal object discovered by D. Rabinowitz with Spacewatch near α = 14^{h}54^{m}, δ = −10^{o}.1 (equinox 2000.0), V = 18.5 on May 21.2 UT moved some 5^{o} to the east and south and faded by 1 mag over the course of 5.5 hr. From an orbit determination by B. G. Marsden, Center for Astrophysics, Rabinowitz appears to have been able to recover the object at V = 20 the following night, but only a single detection could then be made, after which Marsden determined the following improved orbital elements: Epoch = 1993 Aug. 1.0 TT, T = 1993 Apr. 4.60 TT, ω = 261^{o}.30, Ω = 239^{o}.63, i = 3^{o}.19 (equinox 2000.0), q = 0.5017 AU, e = 0.7748, a = 2.2272 AU. The object, with record faint absolute magnitude H ∼ 29.0 (cf. 1991 BA, IAUC 5172), was 0.0047 AU from the earth at discovery and had made a record close approach of 0.0010 AU on May 20.9 TT; minimum distance from the moon was 0.0013 AU on May 20.7.

SUPERNOVA 1993J IN NGC 3031 R. W. Tweedy, Steward Observatory; T. J. Balonek, Colgate Univer-

sity; E. Hintz, M. Joner, and L. Joner, Brigham Young University; and L. A. Wells, Kitt Peak National Observatory, report: “UBVRI CCD photometry at the Burrell-Schmidt telescope on Kitt Peak shows that the U magnitude of SN 1993J has brightened steadily from U = 14.39 during May 23-26 to 14.32 during June 8-11. There is an approximately linear decrease in other bands (0.01 mag/day in B; 0.02 mag/day in V, R, and I). On June 11, B = 13.83, V = 12.79, R = 12.23, I = 11.83, with errors < 0.02 (cf. IAUC 5746, 5769). JHK photometry is being obtained at Kitt Peak. A recent personal communication from C. Wheeler, Univeristy of Texas, mentioned that an increase in U might indicate that the remnant is powered by a pulsar. However, x-ray observations are needed to confirm this.”

FG SAGITTAE G. F. Lawrence and R. D. Gehrz, University of Minnesota; and C. E.

Woodward, University of Wyoming, are continuing their infrared monitoring of FG Sge. Magnitudes obtained with the Mt. Lemmon Observatory 1.52-m telecope (+ InSb photometer): May 31.5 UT, K = 5.78 ± 0.04, L′ = 3.48 ± 0.04, M = 2.56 ± 0.05; June 1.5 UT, J = 8.16 ± 0.03, H = 7.21 ± 0.03, K = 5.75 ± 0.02, L′ = 3.61 ± 0.02, M = 2.77 ± 0.05.

1993 June 18 Daniel W. E. Green